OpenStudy (anonymous):

What is the solution of the equation? 1+ sqrt 3x+3= sqrt 7x+2

OpenStudy (anonymous):

sqrtrt covers the 3x+3?

OpenStudy (anonymous):

yes and on the other side, the square root also covers the 7x+2

OpenStudy (anonymous):

x=2, if the square root sign covers the 3x+3 and 7x+2 within its boundaries

OpenStudy (kc_kennylau):

I think he means: \[\sqrt{3x+3}=\sqrt{7x+2}\]

OpenStudy (anonymous):

x=2 is ur answers

OpenStudy (anonymous):

answer* singular

OpenStudy (kc_kennylau):

just square both sides I think?

OpenStudy (anonymous):

@samsterz. I put in the 2 but the equations aren't equal

OpenStudy (anonymous):

the answer is x=2 u must f calcualted wrong

OpenStudy (anonymous):

Oh nvm Yes they are. Thanks :D